# Singular bifurcations in a slow-fast modified Leslie-Gower model with Holling type II functional response, weak Allee effect and a generalist predator

### v1

@@ Abstract.
We study a predator-prey system with a generalist Leslie-Gower predator, a functional Holling type II response, and a weak Allee effect on the prey. The prey’s population often grows much faster than its predator, allowing us to introduce a small time scale parameter $\varepsilon$ that relates the growth rates of both species, giving rise to a slow-fast system. Zhu and Liu (2022) show that, in the case of the weak Allee effect, Hopf singular bifurcation, slow-fast canard cycles, relaxation oscillations, etc., exist. Our main contribution lies in the rigorous analysis of a degenerate scenario organized by a (degenerate) transcritical bifurcation. The key tool employed is the blow-up method that desingularizes the degenerate singularity. In addition, we determine the criticality of the singular Hopf bifurcation using recent intrinsic techniques that do not require a local normal form. The theoretical analysis is complemented by a numerical bifurcation analysis, in which we numerically identify and analytically confirm the existence of a nearby Takens-Bogdanov point.
@@ 1. Introduction
Mathematical ecology relies heavily on the analysis of predator-prey dynamics to forecast species interactions and the evolution, growth, and population distribution patterns of the biological species involved. The earliest predator-prey model based on sound mathematical principles was proposed by Lotka [27] and Volterra [37]. Following their findings, numerous attempts to improve them have been carried out since then. The Leslie-Gower model is one of the efforts to achieve this task, whose main characteristic is that the predator’s growth equation is logistic, which implies that there is competition or self-interference between the predators, see [25]. Currently, the Leslie-Gower model has been studied from mathematical as well as biological viewpoints by many researchers, which is evident from the amount of published works, see for instance, [14], [20], [10], [31], [13] and references therein.
In 1931, the biologist W.C. Allee investigated population fluctuations and found out that a variety of biological phenomena affect species dynamics, such as mate-finding difficulties, social thermoregulation, genetic drift, avoidance of natural enemies, avoidance of predators, the defense of resources and poor nutrition due to low population densities. This phenomenon is now referred to as the Allee effect. Strong Allee effects result in critical population sizes, whereas weak Allee effects do not. This is due to the effect’s link between population density or size to mean individual fitness. The readers who are interested in this topic are encouraged to refer to, e.g., [30], [3], [33].
The functional response, which refers to the amount of prey a predator consumes per unit of time, is a crucial component of mathematical models of predator-prey dynamics. The ability of the prey to flee an attack or the success of the predator’s pursuit are two examples of the variables that it may affect. Numerous population studies have made the assumption that the functional response is of Holling type I, II, or III; see [17], [18]. The predator-prey model with a functional Holling type II response is the subject of this study. This type of response indicates that the predator spends more time hunting for prey at low prey densities and more time handling prey at high prey densities. A functional Holling type II response is expressed mathematically as:
@@
This article revisits the slow-fast system (1) also considered by Zhu and Liu [42], focusing on a previously unaddressed degenerate case that displays intriguing dynamics characterized by the interaction of a Hopf and a degenerate transcritcal singularities. This degenerate case occurs when the slow nullcline intersects the fast nullcline along the $y$-axis, as detailed in Section 2. In the vicinity of this case, we also uncover novel phenomena through numerical simulations that have not been previously reported. Moreover, we use modern techniques to evaluate the criticality of the organizing singularity. In contrast to the findings of Zhu and Liu [42], our analysis identifies the conditions under which degenerate relaxation oscillations persist in a scenario organized by a transcritical (degenerate) bifurcation.
@@ 2. Existence and types of positive equilibria
$\begin{array}[]{l}x=Ku,\qquad y=\dfrac{nK^{2}}{q}v,\qquad d\tau=\dfrac{rnK^{3}}{(nx+c)(u+a)}\textup{d}t,\qquad A=\dfrac{a}{K}<1,\vspace{0.1cm}\\ C=\dfrac{c}{Kn},\qquad S=\dfrac{s}{rK},\qquad Q=\dfrac{rK}{nq},\qquad M=\dfrac{m}{K},\vspace{0.1cm}\\ \end{array}$
so that $(A,M,C,S,Q)\in(0,1)\times(-1,1)\times\mathbb{R}^{3}_{+}$. In this way, we have a system with only 5 free parameters. Substituting the new variables into system (1) and recycling $t$ to denote the rescaled time $\tau$ yields
@@
The following result shows that the dynamics of the system, taking positive initial conditions, is contained in a region of the first quadrant. Moreover, the solutions are bounded, and they do not depend on the value of $M$. This fact is biologically important because of the achievability and finiteness of populations.
@@
and also $Sv(u+C-Qv)<Sv(1+C-Qv)$. Moreover, $u+A<1+A$ implies that $Sv(u+A)(u+C-Qv)<Sv(1+A)(1+C-Qv).$ Thus, from the second equation of (2), one has
@@
  • Consider case $3$ in Theorem 1 with $A$ and $M$ fixed. Let $Q$ be small enough that there is exactly one positive equilibrium point $E_{1}=(\mathcal{U}_{1},\mathcal{V}_{1})$. If $A-M-1<0$ then, as the value of $Q$ is increased, the coordinate $\mathcal{U}_{1}$ decreases. Moreover, for $Q=\frac{1}{A-M+AM}$ we have two cases:
@@ 3.1. Linear stability analysis for the case $C<-AMQ$
Let the system (2) be such that, we are in case 3(a) of Theorem 1  [^1] Where $M<0$, $C<-AMQ$, $A-M-1<0$ and $M-A-AM+\frac{1}{Q}<0$ or $A-M-1>0$ and $M-A-AM+\frac{1}{Q}>0$ or $A-M-1>0$ and $M-A-AM+\frac{1}{Q}<0$.. Then, the system (2) has only one positive interior equilibrium point $E_{1}=(\mathcal{U}_{1},\mathcal{V}_{1})$, which is:
@@ 4. Analysis of the slow-fast model
$\begin{array}[]{l}\dfrac{\textup{d}u}{\textup{d}t}=f(u,v)=\>u(u+C)\big{(}(u+A)(1-u)(u-M)-v\big{)},\vspace{0.2cm}\\ \dfrac{\textup{d}v}{\textup{d}t}=\varepsilon g(u,v)=\>\varepsilon v(u+A)(u-Qv+C).\end{array}$ (16)
@@
$\frac{\textup{d}u}{\textup{d}\tau}\bigg{|}_{\mathcal{M}_{0}^{1}}=\frac{g(u,h(u))}{h^{\prime}(u)}=\frac{h(u)(u+A)(u+C-Qh(u))}{h^{\prime}(u)}$ (26)
@@ 5. Intrinsic determination of the criticality of the singular Hopf bifurcation
By using the inherent properties of the notion of a slow-fast Hopf point, De Maesschalck et al. [4] prove that the criticality of a slow-fast Hopf can be determined without the necessity of having or changing to a normal form. The authors present a single intrinsic formula for an intrinsic determination that verifies if any singular contact point of Hopf type exists and its criticality. Such an intrinsic formulation is beneficial because the critical curve does not require a parametrization. We use such a technique in the following to show that system (16) may undergo a supercritical Hopf bifurcation as the parameter $Q$ passes through $Q_{H}$.
@@ 6. Desingularization of the point $T_{\mathcal{C}}$, in the degenerate case $C=-AMQ$
For reasons that become clear in the blow-up analysis, we introduce the non-degeneracy conditions $A-M+AM\neq 0$ and $A-M+AM\neq\frac{1}{Q}$.
@@ 6.1. Entry chart
$u=r_{1}u_{1},\quad v=r_{1},\quad\varepsilon=r_{1}\epsilon_{1},$ (34)
@@
$\begin{split}\frac{\textup{d}r_{1}}{\textup{d}t_{1}}&=r_{1}\epsilon_{1}(Q-u_{1})(A^{2}M+\mathcal{O}(r_{1})),\\ \frac{\textup{d}u_{1}}{\textup{d}t_{1}}&=AMQ(1-(A+M-AM)u_{1})u_{1}-A^{2}M(Q-u_{1})u_{1}\epsilon_{1}+\mathcal{O}(r_{1}),\\ \frac{\textup{d}\epsilon_{1}}{\textup{d}t_{1}}&=-\epsilon_{1}^{2}(Q-u_{1})(A^{2}M+\mathcal{O}(r_{1})),\end{split}$ (35)
@@
$\begin{split}\frac{\textup{d}u_{1}}{\textup{d}t_{1}}&=AMQ(1-(A+M-AM)u_{1})u_{1}-A^{2}M(Q-u_{1})u_{1}\epsilon_{1},\\ \frac{\textup{d}\epsilon_{1}}{\textup{d}t_{1}}&=-A^{2}M\epsilon_{1}^{2}(Q-u_{1}).\end{split}$ (36)
@@
The origin is a semi-hyperbolic equilibrium point of (36) and has a unique $1$-dimensional, locally attracting, center manifold $\mathcal{M}_{1}$ associated to it. For $\epsilon_{1}\geq 0$, $\mathcal{M}_{1}$ is unique, coincides with the $\epsilon_{1}$-axis, and the flow along it is directed away from the origin.
@@
System (36) has another equilibrium point at $(u_{1},\epsilon_{1})=\left(\frac{1}{A-M+AM},0\right)$ but this is studied in the chart $K_{3}$.
@@
Next, the dynamics of (35) restricted to $\epsilon_{1}=0$ are given by:
@@
Notice that in this chart $\varepsilon=r_{1}\epsilon_{1}$. Therefore, rescaling time, it is possible to readily see from (35) that the slow flow along the $r_{1}$ axis is given by $r_{1}^{\prime}=A^{2}MQ<0$. See Figure 9.
@@
$\begin{split}\Delta_{1}^{\textup{in}}&=\left\{(r_{1},u_{1},\epsilon_{1})\,:\,r_{1}=\delta_{1},\,0\leq u_{1}<\tilde{u}_{1},\,0\leq\epsilon_{1}<\delta_{2}\right\}\\ \Delta_{1}^{\textup{out}}&=\left\{(r_{1},u_{1},\epsilon_{1})\,:\epsilon_{1}=\delta_{2},\,0\leq u_{1}<\tilde{u}_{1},\,0\leq r_{1}<\delta_{1}\right\},\end{split}$ (39)
@@
Refer to caption $\epsilon_{1}$$u_{1}$$r_{1}$$\mathcal{M}_{1}$$\Delta_{1}^{\textup{in}}$$\Delta_{1}^{\textup{out}}$
Figure 9. Following the analysis presented in this section, we provide a sketch of the dynamics of (35) near the origin. The limit dynamics are presented in green and blue (for $\epsilon_{1}=0$ and $r_{1}=0$ respectively), while a sample orbit of (35) is shown in red.
@@ 6.3. Exit Chart
$u=r_{3},\quad v=r_{3}v_{3},\quad\varepsilon=r_{3}\epsilon_{3},$ (47)
@@
$\begin{split}\frac{\textup{d}r_{3}}{\textup{d}t_{3}}&=-AMQr_{3}(A-M+AM-v_{3})\\ &\;+(A-M+AM-AMQ+A^{2}MQ-AM^{2}Q-v_{3})r_{3}^{2}+\mathcal{O}(r_{3}^{3}),\\ \frac{\textup{d}v_{3}}{\textup{d}t_{3}}&=AMQ(A-M+AM-v_{3})v_{3}+A^{2}M(Qv_{3}-1)\epsilon_{3}+\mathcal{O}(r_{3}),\\ \frac{\textup{d}\epsilon_{3}}{\textup{d}t_{3}}&=AMQ\epsilon_{3}(A-M+AM-v_{3})+\mathcal{O}(\epsilon_{3}r_{3}),\end{split}$ (48)
where $t_{3}$ denotes the rescaled time in this chart and $r_{3}\geq 0$, $\epsilon_{3}\geq 0$. First, we consider (48) restricted to $r_{3}=0$, that is:
$\begin{split}\frac{\textup{d}v_{3}}{\textup{d}t_{3}}&=AMQ(A-M+AM-v_{3})v_{3}+A^{2}M(Qv_{3}-1)\epsilon_{3},\\ \frac{\textup{d}\epsilon_{3}}{\textup{d}t_{3}}&=AMQ\epsilon_{3}(A-M+AM-v_{3}).\end{split}$ (49)
@@
The first two statements are straightforward. For the third, standard center manifold computations show that $\mathcal{N}_{3}=\{(v_{3},\epsilon_{3})\in\mathbb{R}^{2}\,|\,v_{3}=g_{3}(\epsilon_{3})\}$ is given by
$g_{3}(\epsilon_{3})=A-M+AM+\frac{A(A-M+AM-\frac{1}{Q})}{(A-M+AM)Q}\epsilon_{3}+\frac{A^{2}(A-M+AM-\frac{1}{Q})}{(A-M+AM)Q^{2}}\epsilon_{3}^{2}+\mathcal{O}(\epsilon_{3}^{3}),$ (51)
@@
$\frac{\textup{d}\epsilon_{3}}{\textup{d}t_{3}}\biggm{|}_{\mathcal{N}_{3}}=-\frac{A^{2}M(A-M+AM-\frac{1}{Q})}{(A-M+AM)}\epsilon_{3}^{2}++\mathcal{O}(\epsilon_{3}^{3}),$ (52)
@@
On the other hand, system (48) restricted to $\epsilon_{3}=0$ is given by:
@@
Let $A-M+AM-\frac{1}{Q}>0$. There is a (non-unique) $2$-dimensional center manifold $\mathcal{P}_{3}$ at the point $p_{3}$. The flow restricted to it is of saddle type, with the direction $r_{3}$ stable, and the direction $\epsilon_{3}$ unstable, as sketched in Figure 11.
@@
The result follows from standard center manifold computations, so we simply report the important steps. Assuming that the center manifold can be given as the graph of $v_{3}=H_{3}(r_{3},\epsilon_{3})$, we find that
$H_{3}=A-M+AM+\frac{-1+(A-M+AM)Q}{(A-M+AM)Q}A\epsilon_{3}+(1-A+M)r+O(r_{3}^{2},\epsilon_{3}^{2},r_{3}\epsilon_{3}).$ (55)
@@
$\begin{split}\frac{\textup{d}r_{3}}{\textup{d}t_{3}}&=\frac{r_{3}(AMQ-r_{3})(A\epsilon_{3}(-1+(A-M+AM)Q)+(A-M+AM)Qr^{2})}{(A-M+AM)Q}\\ \frac{\textup{d}\epsilon_{3}}{\textup{d}t_{3}}&=-\frac{\epsilon_{3}(AMQ-r_{3})(A\epsilon_{3}(-1+(A-M+AM)Q)+(A-M+AM)Qr^{2})}{(A-M+AM)Q}.\end{split}$ (56)
The condition $A-M+AM-\frac{1}{Q}>0$, together with $M<0$ and $A-M+AM>0$, imply that the common term $\frac{A\epsilon_{3}(-1+(A-M+AM)Q)+(A-M+AM)Qr^{2}}{(A-M+AM)Q}$ is positive. The statement hence follows. ∎
@@
$\begin{split}\Delta_{3}^{\textup{in}}&=\left\{(r_{3},v_{3},\epsilon_{3})\,:\,\epsilon_{3}=\delta_{6},\,|v_{3}-h_{3}(\delta_{6})|<\delta_{7},\,0\leq r_{3}<\tilde{r}_{3}\right\},\\ \Delta_{c}^{\textup{in}}&=\left\{(r_{3},v_{3},\epsilon_{3})\,:\,r_{3}=\rho_{3},\,|v_{3}-\ell_{3}|<e^{-c/\varepsilon},\,0\leq\epsilon_{3}<\delta_{8}\right\}\\ \Delta_{3}^{\textup{out}}&=\left\{(r_{3},v_{3},\epsilon_{3})\,:\,r_{3}=\delta_{9},\,|v_{3}|<\delta_{10},\,0\leq\epsilon_{3}<\tilde{\epsilon}_{3}\right\},\end{split}$ (57)
where all the introduced constants are sufficiently small and positive, further satisfying $\tilde{\epsilon}_{3}<\delta_{6}$, $\tilde{r}_{3}<\delta_{9}$, $\delta_{8}<\delta_{6}$. Let the maps $\Pi_{3}:\Delta_{3}^{\textup{in}}\to\Delta_{3}^{\textup{out}}$ and $\Pi_{3}^{c}:\Delta_{c}^{\textup{in}}\to\Delta_{3}^{\textup{out}}$ be defined by the flow of (48). The analysis presented above suffices to show the following:

### v2

# Singular bifurcations in a modified Leslie-Gower model

@@ Abstract.
We study a predator-prey system with a generalist Leslie-Gower predator, a functional Holling type II response, and a weak Allee effect on the prey. The prey’s population often grows much faster than its predator, allowing us to introduce a small time scale parameter $\varepsilon$ that relates the growth rates of both species, giving rise to a slow-fast system. Zhu and Liu (2022) show that, in the case of the weak Allee effect, Hopf singular bifurcation, slow-fast canard cycles, relaxation oscillations, etc. Our main contribution lies in the rigorous analysis of a degenerate scenario organized by a (degenerate) transcritical bifurcation. The key tool employed is the blow-up method that desingularizes the degenerate singularity. In addition, we determine the criticality of the singular Hopf bifurcation using recent intrinsic techniques that do not require a local normal form. The theoretical analysis is complemented by a numerical bifurcation analysis, in which we numerically identify and analytically confirm the existence of a nearby Takens-Bogdanov point.
@@ 1. Introduction
Mathematical ecology relies heavily on the analysis of predator-prey dynamics to forecast species interactions and the evolution, growth, and population distribution patterns of the biological species involved. The earliest predator-prey model based on sound mathematical principles was proposed by Lotka [27] and Volterra [37]. Following their findings, numerous attempts to improve them have been carried out since then. The Leslie-Gower model is one of the efforts to achieve this task, whose main characteristic is that the predator’s growth equation is logistic, which implies that there is competition or self-interference between the predators, see [25]. Currently, the Leslie-Gower model has been studied from mathematical as well as biological viewpoints by many researchers, which is evident from the amount of published works, see for instance, [10, 13, 14, 20, 31] and references therein.
In 1931, the biologist W.C. Allee investigated population fluctuations and found out that a variety of biological phenomena affect species dynamics, such as mate-finding difficulties, social thermoregulation, genetic drift, avoidance of natural enemies, avoidance of predators, the defense of resources and poor nutrition due to low population densities. This phenomenon is now referred to as the Allee effect. Strong Allee effects result in critical population sizes, whereas weak Allee effects do not. This is due to the effect’s link between population density or size to mean individual fitness. The readers who are interested in this topic are encouraged to refer to, e.g., [3, 30, 33].
The functional response, which refers to the amount of prey a predator consumes per unit of time, is a crucial component of mathematical models of predator-prey dynamics. The ability of the prey to flee an attack or the success of the predator’s pursuit are two examples of the variables that it may affect. Numerous population studies have made the assumption that the functional response is of Holling type I, II, or III; see [17, 18]. The predator-prey model with a functional Holling type II response is the subject of this study. This type of response indicates that the predator spends more time hunting for prey at low prey densities and more time handling prey at high prey densities. A functional Holling type II response is expressed mathematically as:
@@
This article revisits the slow-fast system (1) also considered by Zhu and Liu [42], focusing on a previously unaddressed degenerate case that displays intriguing dynamics characterized by the interaction of a Hopf and a degenerate transcritical singularities. This degenerate case occurs when the slow nullcline intersects the fast nullcline along the $y$-axis, as detailed in Section 2. In the vicinity of this case, we also uncover novel phenomena through numerical simulations that have not been previously reported. Moreover, we use modern techniques to evaluate the criticality of the organizing singularity. In contrast to the findings of Zhu and Liu [42], our analysis identifies the conditions under which degenerate relaxation oscillations persist in a scenario organized by a transcritical (degenerate) bifurcation.
@@ 2. Existence and types of positive equilibria
$\begin{array}[]{l}x=Ku,\qquad y=\dfrac{nK^{2}}{q}v,\qquad\textup{d}\tau=\dfrac{rnK^{3}}{(nx+c)(u+a)}\textup{d}t,\qquad A=\dfrac{a}{K}<1,\vspace{0.1cm}\\ C=\dfrac{c}{Kn},\qquad S=\dfrac{s}{rK},\qquad Q=\dfrac{rK}{nq},\qquad M=\dfrac{m}{K},\vspace{0.1cm}\\ \end{array}$
so that $(A,M,C,S,Q)\in(0,1)\times(-1,1)\times\mathbb{R}^{3}_{+}$. In this way, we have a system with only 5 free parameters.
Substituting the new variables into system (1) and recycling $t$ to denote the rescaled time $\tau$ yields
@@
The following result shows that the dynamics of the system, taking positive initial conditions, is contained in a region of the first quadrant. Moreover, the solutions are bounded and do not depend on the value of $M$. This fact is biologically important because of the achievability and finiteness of populations.
@@
and also $Sv(u+C-Qv)<Sv(1+C-Qv)$. Moreover, $u+A<1+A$ implies that
$Sv(u+A)(u+C-Qv)<Sv(1+A)(1+C-Qv).$
Thus, from the second equation of (2), one has
@@
  • Consider case $3$ in Theorem 1 with $A$ and $M$ fixed. Let $Q$ be small enough that there is exactly one positive equilibrium point $E_{1}=(\mathcal{U}_{1},\mathcal{V}_{1})$. If $A-M-1<0$, then as the value of $Q$ is increased, the coordinate $\mathcal{U}_{1}$ decreases. Moreover, for $Q=\frac{1}{A-M+AM}$ we have two cases:
@@ 3.1. Linear stability analysis for the case $C<-AMQ$
Let the system (2) be such that we are in case 3(a) of Theorem 1  [^1] Where $M<0$, $C<-AMQ$, $A-M-1<0$ and $M-A-AM+\frac{1}{Q}<0$ or $A-M-1>0$ and $M-A-AM+\frac{1}{Q}>0$ or $A-M-1>0$ and $M-A-AM+\frac{1}{Q}<0$.. Then, the system (2) has only one positive interior equilibrium point $E_{1}=(\mathcal{U}_{1},\mathcal{V}_{1})$, which is:
@@ 4. Analysis of the slow-fast model
$\begin{array}[]{l}\dfrac{\textup{d}u}{\textup{d}t}=f(u,v)=\>u(u+C)\big{(}(u+A)(1-u)(u-M)-v\big{)},\vspace{0.2cm}\\ \dfrac{\textup{d}v}{\textup{d}t}=\varepsilon g(u,v)=\>\varepsilon v(u+A)(u-Qv+C),\end{array}$ (16)
@@
$\frac{\textup{d}u}{\textup{d}\tau}\bigg{|}_{\mathcal{M}_{0}^{1}}=\frac{g(u,h(u))}{h^{\prime}(u)}=\frac{h(u)(u+A)(u+C-Qh(u))}{h^{\prime}(u)}.$ (26)
@@ 5. Intrinsic determination of the criticality of the singular Hopf bifurcation
By using the inherent properties of the notion of a slow-fast Hopf point, De Maesschalck and co-authors [4] prove that the criticality of a slow-fast Hopf can be determined without the necessity of having or changing to a normal form. The authors present a single intrinsic formula for an intrinsic determination that verifies if any singular contact point of Hopf type exists and its criticality. Such an intrinsic formulation is beneficial because the critical curve does not require a parametrization. We use such a technique in the following to show that system (16) may undergo a supercritical Hopf bifurcation as the parameter $Q$ passes through $Q_{H}$.
@@ 6. Desingularization of the point $T_{\mathcal{C}}$, in the degenerate case $C=-AMQ$
For reasons that become clear in the blow-up analysis, see in particular Propositions 6, 7, 8, 9, 10 and Theorem 4, we introduce the non-degeneracy conditions $A-M+AM\neq 0$ and $A-M+AM\neq\frac{1}{Q}$.
@@ 6.1. Entry chart
$u=r_{1}u_{1},\quad v=r_{1},\quad\varepsilon=r_{1}\varepsilon_{1},$ (34)
@@
$\begin{split}\frac{\textup{d}r_{1}}{\textup{d}t_{1}}&=r_{1}\varepsilon_{1}(Q-u_{1})(A^{2}M+\mathcal{O}(r_{1})),\\ \frac{\textup{d}u_{1}}{\textup{d}t_{1}}&=AMQ(1-(A+M-AM)u_{1})u_{1}-A^{2}M(Q-u_{1})u_{1}\varepsilon_{1}+\mathcal{O}(r_{1}),\\ \frac{\textup{d}\varepsilon_{1}}{\textup{d}t_{1}}&=-\varepsilon_{1}^{2}(Q-u_{1})(A^{2}M+\mathcal{O}(r_{1})),\end{split}$ (35)
@@
$\begin{split}\frac{\textup{d}u_{1}}{\textup{d}t_{1}}&=AMQ(1-(A+M-AM)u_{1})u_{1}-A^{2}M(Q-u_{1})u_{1}\varepsilon_{1},\\ \frac{\textup{d}\varepsilon_{1}}{\textup{d}t_{1}}&=-A^{2}M\varepsilon_{1}^{2}(Q-u_{1}).\end{split}$ (36)
@@
The origin is a semi-hyperbolic equilibrium point of (36) and has a unique $1$-dimensional, locally attracting, center manifold $\mathcal{M}_{1}$ associated to it. For $\varepsilon_{1}\geq 0$, $\mathcal{M}_{1}$ is unique, coincides with the $\varepsilon_{1}$-axis, and the flow along it is directed away from the origin.
@@
System (36) has another equilibrium point at $(u_{1},\varepsilon_{1})=\left(\frac{1}{A-M+AM},0\right)$ but this is studied in the chart $K_{3}$.
@@
Next, the dynamics of (35) restricted to $\varepsilon_{1}=0$ are given by:
@@
Notice that in this chart $\varepsilon=r_{1}\varepsilon_{1}$. Therefore, rescaling time, it is possible to readily see from (35) that the slow flow along the $r_{1}$ axis is given by $r_{1}^{\prime}=A^{2}MQ<0$. See Figure 9.
@@
$\begin{split}\Delta_{1}^{\textup{in}}&=\left\{(r_{1},u_{1},\varepsilon_{1})\,:\,r_{1}=\delta_{1},\,0\leq u_{1}<\tilde{u}_{1},\,0\leq\varepsilon_{1}<\delta_{2}\right\}\\ \Delta_{1}^{\textup{out}}&=\left\{(r_{1},u_{1},\varepsilon_{1})\,:\varepsilon_{1}=\delta_{2},\,0\leq u_{1}<\tilde{u}_{1},\,0\leq r_{1}<\delta_{1}\right\},\end{split}$ (39)
@@
Refer to caption $\varepsilon_{1}$$u_{1}$$r_{1}$$\mathcal{M}_{1}$$\Delta_{1}^{\textup{in}}$$\Delta_{1}^{\textup{out}}$
Figure 9. Following the analysis presented in this section, we provide a sketch of the dynamics of (35) near the origin. The limit dynamics are presented in green and blue (for $\varepsilon_{1}=0$ and $r_{1}=0$ respectively), while a sample orbit of (35) is shown in red.
@@ 6.3. Exit chart
$u=r_{3},\quad v=r_{3}v_{3},\quad\varepsilon=r_{3}\varepsilon_{3},$ (47)
@@
$\begin{split}\frac{\textup{d}r_{3}}{\textup{d}t_{3}}&=-AMQr_{3}(A-M+AM-v_{3})\\ &\;+(A-M+AM-AMQ+A^{2}MQ-AM^{2}Q-v_{3})r_{3}^{2}+\mathcal{O}(r_{3}^{3}),\\ \frac{\textup{d}v_{3}}{\textup{d}t_{3}}&=AMQ(A-M+AM-v_{3})v_{3}+A^{2}M(Qv_{3}-1)\varepsilon_{3}+\mathcal{O}(r_{3}),\\ \frac{\textup{d}\varepsilon_{3}}{\textup{d}t_{3}}&=AMQ\varepsilon_{3}(A-M+AM-v_{3})+\mathcal{O}(\varepsilon_{3}r_{3}),\end{split}$ (48)
where $t_{3}$ denotes the rescaled time in this chart and $r_{3}\geq 0$, $\varepsilon_{3}\geq 0$. First, we consider (48) restricted to $r_{3}=0$, that is:
$\begin{split}\frac{\textup{d}v_{3}}{\textup{d}t_{3}}&=AMQ(A-M+AM-v_{3})v_{3}+A^{2}M(Qv_{3}-1)\varepsilon_{3},\\ \frac{\textup{d}\varepsilon_{3}}{\textup{d}t_{3}}&=AMQ\varepsilon_{3}(A-M+AM-v_{3}).\end{split}$ (49)
@@
The first two statements are straightforward. For the third, standard center manifold computations show that $\mathcal{N}_{3}=\{(v_{3},\varepsilon_{3})\in\mathbb{R}^{2}\,|\,v_{3}=g_{3}(\varepsilon_{3})\}$ is given by
$g_{3}(\varepsilon_{3})=A-M+AM+\frac{A(A-M+AM-\frac{1}{Q})}{(A-M+AM)Q}\varepsilon_{3}+\frac{A^{2}(A-M+AM-\frac{1}{Q})}{(A-M+AM)Q^{2}}\varepsilon_{3}^{2}+\mathcal{O}(\varepsilon_{3}^{3}),$ (51)
@@
$\frac{\textup{d}\varepsilon_{3}}{\textup{d}t_{3}}\biggm{|}_{\mathcal{N}_{3}}=-\frac{A^{2}M(A-M+AM-\frac{1}{Q})}{(A-M+AM)}\varepsilon_{3}^{2}+\mathcal{O}(\varepsilon_{3}^{3}),$ (52)
@@
On the other hand, system (48) restricted to $\varepsilon_{3}=0$ is given by:
@@
Let $A-M+AM-\frac{1}{Q}>0$. There is a (not necessarily unique) $2$-dimensional center manifold $\mathcal{P}_{3}$ at the point $p_{3}$. The flow restricted to it is of saddle type, with the direction $r_{3}$ stable and the direction $\varepsilon_{3}$ unstable, as sketched in Figure 11.
@@
The result follows from standard center manifold computations, so we simply report the important steps. Assuming that the center manifold can be given as the graph of $v_{3}=H_{3}(r_{3},\varepsilon_{3})$, we find
$H_{3}=A-M+AM+\frac{-1+(A-M+AM)Q}{(A-M+AM)Q}A\varepsilon_{3}+(1-A+M)r+\mathcal{O}(r_{3}^{2},\varepsilon_{3}^{2},r_{3}\varepsilon_{3}).$ (55)
@@
$\begin{split}\frac{\textup{d}r_{3}}{\textup{d}t_{3}}&=\frac{r_{3}(AMQ-r_{3})(A\varepsilon_{3}(-1+(A-M+AM)Q)+(A-M+AM)Qr^{2})}{(A-M+AM)Q},\\ \frac{\textup{d}\varepsilon_{3}}{\textup{d}t_{3}}&=-\frac{\varepsilon_{3}(AMQ-r_{3})(A\varepsilon_{3}(-1+(A-M+AM)Q)+(A-M+AM)Qr^{2})}{(A-M+AM)Q}.\end{split}$ (56)
The condition $A-M+AM-\frac{1}{Q}>0$, together with $M<0$ and $A-M+AM>0$, imply that the common term $\frac{A\varepsilon_{3}(-1+(A-M+AM)Q)+(A-M+AM)Qr^{2}}{(A-M+AM)Q}$ is positive. In addition, recalling Proposition 8, we know that $p_{3}$ is a non-hyperbolic source, suggesting the non-uniqueness of the center manifold. The statement therefore follows. ∎
@@
$\begin{split}\Delta_{3}^{\textup{in}}&=\left\{(r_{3},v_{3},\varepsilon_{3})\,:\,\varepsilon_{3}=\delta_{6},\,|v_{3}-h_{3}(\delta_{6})|<\delta_{7},\,0\leq r_{3}<\tilde{r}_{3}\right\},\\ \Delta_{c}^{\textup{in}}&=\left\{(r_{3},v_{3},\varepsilon_{3})\,:\,r_{3}=\rho_{3},\,|v_{3}-\ell_{3}|<e^{-c/\varepsilon},\,0\leq\varepsilon_{3}<\delta_{8}\right\}\\ \Delta_{3}^{\textup{out}}&=\left\{(r_{3},v_{3},\varepsilon_{3})\,:\,r_{3}=\delta_{9},\,|v_{3}|<\delta_{10},\,0\leq\varepsilon_{3}<\tilde{\varepsilon}_{3}\right\},\end{split}$ (57)
where all the introduced constants are sufficiently small and positive, further satisfying $\tilde{\varepsilon}_{3}<\delta_{6}$, $\tilde{r}_{3}<\delta_{9}$, $\delta_{8}<\delta_{6}$. Let the maps $\Pi_{3}:\Delta_{3}^{\textup{in}}\to\Delta_{3}^{\textup{out}}$ and $\Pi_{3}^{c}:\Delta_{c}^{\textup{in}}\to\Delta_{3}^{\textup{out}}$ be defined by the flow of (48). The analysis presented above suffices to show the following: